# A Survey on Recommendation Unlearning: Fundamentals, Taxonomy, Evaluation, and Open Questions

### v1

@@ Abstract
Index Terms:
Recommender Systems, Machine Unlearning, Recommendation Unlearning.
@@ I Introduction
Over the past few years, rapid advancements have propelled machine learning, particularly deep learning, to unprecedented heights [45, 51, 31, 103, 36, 83, 11, 80]. Innovations in algorithms, architectures, and computing power have enabled models to achieve remarkable performance across diverse tasks, revolutionizing industries and reshaping societal norms [92, 93, 44, 9, 111, 101, 23, 1, 22]. Concurrently, exponential growth in data usage for training machine learning models has brought forth pressing concerns, particularly about privacy and security [95, 25, 120]. As more and more personal, sensitive, and potentially harmful data is leveraged for model training, concerns about how this data is collected, stored, and used have intensified. Individuals now face increasing risks to their privacy, with potential misuse or unauthorized access to personal information becoming more prevalent. In this context, regulatory measures such as the General Data Protection Regulation [102], the California Consumer Privacy Act [79], and the Delete Act [40] now empower individuals to exercise their right to be forgotten, demanding the removal of personal data used during model construction.
@@
The challenge, however, extends beyond the simple deletion of data. Machine learning models have been shown to exhibit the ability to memorize or retain information from their training data [43, 37]. This means that even if specific data points are removed from a dataset, traces of that data might persist within the model itself, potentially leading to privacy breaches or non-compliance with regulatory requests. This issue has spurred the development of machine unlearning, i.e., a novel field of research aimed at effectively forgetting specific data that was used during model training [10, 106, 73, 91, 70, 54, 71, 64]. While machine unlearning was initially driven by the need to comply with emerging regulations, its scope has expanded beyond legal compliance, revealing significant potential for improving model quality and performance. Unlearning has proven to be a valuable tool not only for addressing privacy concerns but also for removing data that is inaccurate, toxic, biased, or poisoned [53, 76, 114]. By proactively identifying and eliminating such harmful data, machine unlearning can contribute to the creation of more reliable, fair, and robust machine learning models.
@@
Recommender systems are designed to predict user preferences by analyzing historical interactions, which typically include actions such as clicks, purchases, ratings, and likes. These interactions serve as a foundation for making personalized recommendations and are central to the functioning of many modern digital platforms, from e-commerce websites to streaming services. As machine learning models become more deeply integrated into recommender systems [34, 108, 17, 33, 99, 97, 2, 61], the need for recommendation unlearning has gained significant attention as a critical research area [15]. There are two primary factors that drive the growing necessity of recommendation unlearning. Firstly, historical interactions tend to be intimate and closely tied to user privacy [124, 67, 62, 68, 98], thus carrying a higher likelihood of unlearning requests. Secondly, the precision of recommendation models critically depends on the quality of training data [87], mandating rigorous cleansing regimes to remove dirty data that affects the recommendation performance. As shown in Figure 1, from a practical standpoint, recommendation unlearning involves withdrawal actions within recommendation platforms. This can include activities such as revoking or modifying past likes, ratings, or other forms of feedback, reflecting the dynamic nature of user engagement with platforms. Users may update their preferences, delete or modify past interactions, or request that certain types of recommendations be forgotten. The ability to unlearn data from a model ensures that it remains up-to-date, accurate, and in compliance with user requests and regulatory requirements.
@@
Figure 1: Illustration of applications scenarios of recommendation unlearning (e.g., social media, entertainment, and e-commerce).
@@
However, traditional unlearning methods, primarily tailored for image classification tasks, are poorly suited to the complex requirements of recommendation unlearning [56]. This misalignment arises from the fundamentally different nature of recommender systems, which rely heavily on collaborative interactions between users and items to provide personalized and relevant suggestions. In contrast, image classification tasks typically involve a more straightforward relationship between inputs and labels, without the intricate dependencies that characterize recommender systems. Traditional unlearning methods, if applied indiscriminately in recommender systems, risk disrupting these delicate collaborative relationships, leading to the loss of crucial patterns that underpin accurate recommendations [15, 58, 57]. As a result, removing data from the model without careful consideration can degrade the model’s performance, rendering it less effective at making personalized predictions and reducing the overall quality of the recommendations. Furthermore, the vast parameters of recommendation models, particularly user and item embeddings, exacerbate the inefficiency issue of traditional unlearning [58]. Given these unique challenges, it is clear that recommendation unlearning cannot rely on traditional unlearning techniques. Instead, new and specialized methods are required.
@@
Given the growing significance of recommendation unlearning, merely one existing survey have reviewed the existing techniques employed in this area [86]. While such a survey provide valuable insights into the methods currently in use, they often focus narrowly on the technical aspects without offering a holistic view of the emerging challenges and opportunities in the domain. In light of this gap, the aim of this survey is to provide a more comprehensive and in-depth review of the latest advancements in recommendation unlearning, while also exploring potential research directions that could shape the future of this field. This survey not only evaluates existing methodologies but also synthesizes key insights into the design principles and challenges that researchers need to address going forward. We summarize the main contribution of our survey as follows:
@@
  • Foundational Introduction (Section II): Beginning with foundational concepts of recommendation unlearning, we establish a clear understanding of recommendation unlearning target, unlearning workflow, and design principles that guide the development of unlearning techniques.
@@ II-A1 Machine Unlearning
Machine unlearning aims to eliminate the memory of specific training from a given model. Given a dataset $\mathcal{D}$ and a learning algorithm $A(\cdot)$, a model $\theta_{0}=A(\mathcal{D})$ can be learned (i.e., the original model before unlearning). An unlearning task typically specifies a target that needs to be unlearned. In earlier literature, the unlearning targets were generally confined to the training data, as the model’s memory is learned directly from the training data [12, 10]. We refer to these traditional unlearning tasks as input unlearning, meaning they focus on unlearning the input data used to train the model. A more detailed discussion of different unlearning targets is provided in Section II-B. For clarity and simplicity, we define the unlearning target here as a subset of the training data (i.e., the forget set $\mathcal{D}_{f}\in\mathcal{D}$). The remaining training data is therefore defined as the retrain set, where $\mathcal{D}_{f}\cup\mathcal{D}_{r}=\mathcal{D}$.
As shown in Figure 2, a straightforward method of unlearning is to retrain the model from scratch using the retain set (i.e., $\theta^{*}=A(\mathcal{D}_{r})$) However, with the increasing size of both data and models in modern applications, retraining from scratch becomes computationally prohibitive and impractical. This challenge has, in turn, driven the development of more efficient methods for machine unlearning. The retrained model $\theta^{*}$ serves as the ground truth for unlearning. Therefore, the goal of unlearning is to design an unlearning algorithm $U(\cdot)$ such that
@@
[…]
where the input of $U(\cdot)$ may be reduced under different conditions (e.g., when the retain set is not available), $Cost(\cdot)$ denotes the computational overhead, and $\doteq$ denotes the degree of unlearning completeness, i.e., to what extent the target has been completely unlearned. This is also referred to as the effectiveness or efficacy of unlearning in the literature. The discussion of evaluation metrics of unlearning completeness will be provided in Section II-D. For clarity, in this paper, we refer to $\theta$ as the original model, $\theta_{u}=U(A,\mathcal{D}_{f},\mathcal{D}_{r},\theta_{0})$ as the unlearned model, and $\theta^{*}=A(\mathcal{D}_{r})$ as the retrained model.
@@ II-A2 Recommender systems
Recommender systems leverage historical interactions to predict user preference. Most research focuses on the rating prediction task, where users rate various items [96]. Recommendation models leverage observed (i.e., historical) ratings to predict unobserved ones. Collaborative filtering is the foundational approach in modern recommender systems, based on the assumption that users tend to favor similar items, and that an item is likely to be favored by similar users, creating a collaborative effect [49]. Under this approach, matrix factorization-based models have become widely used in recommender systems, both in academia and industry [75, 38]. Existing research on recommendation unlearning also primarily focuses on matrix factorization-based models. Generally, the core idea behind matrix factorization-based models is to learn a user embedding matrix and an item embedding matrix. The predicted ratings are then obtained through the dot product (or other combination methods) of these two matrices. Subsequent research has leveraged deep learning techniques [116, 34], graph learning approaches [33, 81, 125], and Large Language Models (LLMs) [2, 61] to further enhance recommendation performance.
@@ II-A3 Recommendation Unlearning
  • Last but not least, recommendation models are highly sensitive to the quality of the training data [87], creating a need for systems to unlearn problematic data proactively.
@@
However, traditional machine unlearning methods cannot be indiscriminately applied to recommender systems, due to the unique challenges posed by collaborative interactions and parameters of recommendation models. Firstly, as shown in Figure 3, due to the collaborative effect of recommendation data, unlearning the forget set may inadvertently affect the associated data, disrupting its memory. This causes a breakdown in the quality of recommendations and potentially leads to a diminished user experience. Conversely, the retain set can leverage the collaborative effects to re-remember the forget set, thereby causing non-compliance with the unlearning request. Secondly, the user and item embeddings, which are crucial parameters in recommendation models, capture the features of users and items. As a result, existing recommendation unlearning methods typically focus on manipulating these embeddings. However, because the user and item embeddings form high-dimensional and dense matrices, and the number of users and items is massive, applying traditional unlearning methods to these embeddings introduces significant computational overhead, which makes recommendation unlearning even more challenging.
II-B Unlearning Targets
@@
In addition to the training data, data that does not participate in training may also need to be unlearned in recommender systems. This is because adversaries can potentially infer private information from a trained model, even if that information was never explicitly included in the training data. This type of information, referred to as attributes, is implicitly learned by the model during training. Such attacks are known as attribute inference attacks [27, 123, 41]. The task of unlearning these attributes is called attribute unlearning, which serves as a defense mechanism against attribute inference attacks [26, 59, 18]. As shown in Figure 2, the unlearning target of attribute unlearning is the latent user attributes that are not part of the training data, e.g., gender, age, and race.
@@ II-C Unlearning Workflow
II-D Design Principles
The goal of unlearning is not merely to eliminate the memory of the target being unlearned. It encompasses broader goals. Generally, there are three key design principles for unlearning methods, which are also applicable to recommendation scenarios [15, 57, 56].
@@ II-D Design Principles
  • Algorithmic perspective: Only retraining from scratch (i.e., unlearning from the algorithmic level) satisfies the definition of complete unlearning in this perspective. Thus, this definition of completeness can only be self-evaluated algorithmically or verified by providing training checkpoints [42, 100, 109]. Exact unlearning methods adhere strictly to this definition, designing efficient strategies to achieve retraining. All other unlearning methods are classified as approximate unlearning [78].
@@
The specific examples of evaluation metrics are introduced in Section IV-C.
@@ III-A Input Unlearning
Based on the types of models that can be applied, existing input unlearning methods can be categorized into two main approaches: model-agnostic and model-specific. The term model-agnostic refers to methods that can be applied regardless of the model structure, while model-specific refers to methods designed for a particular type of model. We also review the unlearning methods for recommendation tasks in specific scenarios, such as federated learning and LLMs, where the recommendation model differs from traditional recommender systems.
@@ III-A1 Model-agnostic Methods
Model-agnostic unlearning methods for recommendation tasks are largely inspired by traditional machine unlearning techniques used in classification tasks. Based on the definition of unlearning completeness, we further classify the model-agnostic approach into exact unlearning and approximate unlearning [78].
Figure 7: Different designs for exact unlearning in recommender systems, where $\mathcal{D}_{i}$ represents the data subset. The black and blue arrows depict the workflows of learning and unlearning respectively.
@@
As mentioned in Section II-D, exact unlearning follows a strict definition of unlearning completeness, achieving it at the algorithmic level. Inspired by SISA [10], exact recommendation unlearning methods predominantly adopt the ensemble retraining framework. As shown in Figure 7, this framework divides the original dataset into multiple subsets, trains a sub-model on each subset, and aggregates all sub-models into the final model, similar to an ensemble learning system. Note that, to ensure algorithmic unlearning completeness, the design of sub-models is typically identical to that of the original model, including the model structure, hyperparameter settings, and other factors. This framework allows for more efficient retraining during unlearning. When an unlearning request is submitted, the ensemble retraining framework only needs to retrain the subset containing the unlearning targets, avoiding the need to retrain the entire dataset from scratch, thereby enhancing unlearning efficiency.
@@
Li et al. [55] directly apply SISA to recommendation models in intelligence education. Their approach enhances the personalization and accuracy of educational recommendations by selectively forgetting the data inputs for each user.
@@
Building on the design of SISA, Chen et al. propose RecEraser, which introduces two key modifications tailored for recommendation tasks [15]. First, RecEraser incorporates a balanced clustering module for dataset division, grouping similar users or items into the same subset to preserve the collaborative effects within the data, in contrast to the random division used in SISA. Second, RecEraser adds an attention network to learn the weights for the weighted aggregation of sub-models. This adaptive weighted aggregation, compared to the average weighting or majority voting in SISA, further enhances recommendation performance for the ensemble retraining framework.
@@
Due to the collaborative effect of recommendation data, there is a significant trade-off between unlearning efficiency and model utility. Specifically, increasing the number of dataset divisions can enhance unlearning efficiency, but this also disrupts the collaboration among data, which in turn reduces model utility. To address this issue, Li et al. propose UltraRE, a lightweight modification of RecEraser [56]. UltraRE introduces a new balanced clustering algorithm based on optimal transport theory, which improves both efficiency and clustering performance simultaneously. Additionally, UltraRE simplifies the attention network used during aggregation, replacing it with a logistic regression model to further enhance efficiency.
@@
To further enhance model utility, LASER adopts sequential training during aggregation, rather than parallel training [57]. As shown in Figure 7, sequential training involves training one model on a data subset sequentially. This approach helps mitigate the negative impact of dataset division on collaboration. LASER introduces the concept of curriculum learning to optimize the training sequence of data subsets, thereby improving model utility. However, sequential training also reduces unlearning efficiency. To address this issue, LASER introduces early stopping and parameter manipulation.
However, there are several drawbacks of exact unlearning.
@@
  • The efficiency gains from dataset division are limited and incremental, whereas approximate unlearning methods can often provide efficiency improvements that are orders of magnitude better. Additionally, as noted by [56], exact unlearning introduces a trade-off, which is particularly significant in recommendation tasks.
  • The performance of exact unlearning is suboptimal in practice [16]. Although exact unlearning achieves perfect completeness in theory, its empirical performance is unsatisfactory, often worse than approximate unlearning. In real-world scenarios, users prioritize good performance over theoretical guarantees.
Figure 8: Different designs for approximate unlearning in recommender systems, i.e., reverse unlearning and active unlearning.
@@
Reverse Unlearning. Deep learning models are typically trained using gradient descent-based optimization. An intuitive approach for unlearning is to add back the gradient of the target data that was previously subtracted, thereby mitigating its influence on the model and achieving the goal of unlearning from a parametric perspective. Formally, let us assume that $z_{i}\in\mathcal{D}$ is a data record from the training set, and the recommendation loss function is $\ell(z_{i},\theta)$. We further assume that the loss function is twice-differentiable and strictly convex, which holds true for the majority of recommendation loss functions. Following the empirical risk minimization framework, the optimal model parameter is given by
[…]
@@
[…]
The early study in traditional machine unlearning explores directly adding back the first-order gradient [28]. However, due to the stochastic nature and randomness inherent in deep learning, the first-order gradient is not always accurate for estimating the influence of the target data. Consequently, subsequent research has focused on more precisely estimating the influence and directly subtracting this influence from the original model [90]. This process is simply formulated as:

### v2

@@ I Introduction
Over the past few years, rapid advancements have propelled machine learning, particularly deep learning, to unprecedented heights [kingma2014adam, lecun2015deep, he2016deep, vaswani2017attention, ho2020denoising, radford2019language, NEURIPS2020_1457c0d6, openai2023gpt]. Innovations in algorithms, architectures, and computing power have enabled models to achieve remarkable performance across diverse tasks, revolutionizing industries and reshaping societal norms [silver2016mastering, silver2017mastering, jumper2021highly, biderman2023pythia, wu2023bloomberggpt, touvron2023llama, dubey2024llama, abramson2024accurate, dettmers2024qlora]. Concurrently, exponential growth in data usage for training machine learning models has brought forth pressing concerns, particularly about privacy and security [stahl2018ethics, eshete2021making, zhang2024instruction]. As more and more personal, sensitive, and potentially harmful data is leveraged for model training, concerns about how this data is collected, stored, and used have intensified. Individuals now face increasing risks to their privacy, with potential misuse or unauthorized access to personal information becoming more prevalent. In this context, regulatory measures such as the General Data Protection Regulation [euro2018gdpr], the California Consumer Privacy Act [cal2018ccpa], and the Delete Act [cal2023del] now empower individuals to exercise their right to be forgotten, demanding the removal of personal data used during model construction.
@@
The challenge, however, extends beyond the simple deletion of data. Machine learning models have been shown to exhibit the ability to memorize or retain information from their training data [jia2018attriguard, hu2022membership]. This means that even if specific data points are removed from a dataset, traces of that data might persist within the model itself, potentially leading to privacy breaches or non-compliance with regulatory requests. This issue has spurred the development of machine unlearning, i.e., a novel field of research aimed at effectively forgetting specific data that was used during model training [bourtoule2021machine, wang2024machine, liu2024threats, shaik2024exploring, liu2024machine, li2024machine, liu2024survey, liu2024survey2]. While machine unlearning was initially driven by the need to comply with emerging regulations, its scope has expanded beyond legal compliance, revealing significant potential for improving model quality and performance. Unlearning has proven to be a valuable tool not only for addressing privacy concerns but also for removing data that is inaccurate, toxic, biased, or poisoned [li2016data, mehrabi2021survey, xing2024efuf]. By proactively identifying and eliminating such harmful data, machine unlearning can contribute to the creation of more reliable, fair, and robust machine learning models.
@@
Recommender systems are designed to predict user preferences by analyzing historical interactions, which typically include actions such as clicks, purchases, ratings, and likes. These interactions serve as a foundation for making personalized recommendations and are central to the functioning of many modern digital platforms, from e-commerce websites to streaming services. As machine learning models become more deeply integrated into recommender systems [he2017neural, wang2019neural, chen2020efficient, he2020lightgcn, tan20224sdrug, su2023personalized, acharya2023llm, lin2024data], the need for recommendation unlearning has gained significant attention as a critical research area [chen2022recommendation]. There are two primary factors that drive the growing necessity of recommendation unlearning. Firstly, historical interactions tend to be intimate and closely tied to user privacy [zheng2022matrix, liu2023differentially, ijcai2024p0238, liu2024reducing, su2024revisit], thus carrying a higher likelihood of unlearning requests. Secondly, the precision of recommendation models critically depends on the quality of training data [schafer2007collaborative], mandating rigorous cleansing regimes to remove dirty data that affects the recommendation performance. As shown in Figure 1, from a practical standpoint, recommendation unlearning involves withdrawal actions within recommendation platforms. This can include activities such as revoking or modifying past likes, ratings, or other forms of feedback, reflecting the dynamic nature of user engagement with platforms. Users may update their preferences, delete or modify past interactions, or request that certain types of recommendations be forgotten. The ability to unlearn data from a model ensures that it remains up-to-date, accurate, and in compliance with user requests and regulatory requirements.
@@
Figure 1: Illustration of application scenarios of recommendation unlearning (e.g., social media, entertainment, and e-commerce).
@@
However, traditional unlearning methods, primarily tailored for image classification tasks, are poorly suited to the complex requirements of recommendation unlearning [li2023ultrare]. This misalignment arises from the fundamentally different nature of recommender systems, which rely heavily on collaborative interactions between users and items to provide personalized and relevant suggestions. In contrast, image classification tasks typically involve a more straightforward relationship between images and labels, without the collaborative dependencies that characterize recommender systems. Traditional unlearning methods, if applied indiscriminately in recommender systems, risk disrupting these delicate collaborative relationships, leading to the loss of crucial patterns that underpin accurate recommendations [chen2022recommendation, li2023selective, li2024making]. Furthermore, the vast parameters of recommendation models, particularly user and item embeddings, exacerbate the inefficiency issue of traditional unlearning [li2023selective]. Thus, the collaborative dependencies can create additional challenges for recommendation unlearning in terms of target identification, method design, and unlearning evaluation. Given these unique challenges, it is clear that recommendation unlearning cannot rely on traditional unlearning techniques. Instead, new and specialized methods are required.
@@
Given the growing significance of recommendation unlearning, only one existing survey has reviewed the existing techniques employed in this area [sachdeva2024machine]. While such a survey provides valuable insights into current methods, it focuses narrowly on the technical aspects, failing to establish a systematic taxonomy of these methods. Moreover, it provides limited information on evaluation methodologies, emerging challenges, and future opportunities in the domain. This narrow scope leaves significant gaps in our comprehensive understanding of recommendation unlearning. In light of this gap, the aim of this survey is to deliver a more comprehensive and in-depth review of i) technical approaches along with their taxonomy and design principles, ii) evaluation methodologies and widely-used metrics, iii) summarization of existing methods and research trends, and iv) open questions and future research directions. We summarize the main contribution of our survey as follows:
@@
  • Foundational Introduction (Section II): Beginning with foundational concepts of recommendation unlearning, we establish a clear understanding of the recommendation unlearning target, unlearning workflow, and design principles that guide the development of unlearning techniques.
@@ II-A1 Machine Unlearning
Machine unlearning aims to eliminate the memory of specific training from a given model. Given a dataset $\mathcal{D}$ and a learning algorithm $A(\cdot)$, a model $\theta_{0}=A(\mathcal{D})$ can be learned (i.e., the original model before unlearning). An unlearning task typically specifies a target that needs to be unlearned. In earlier literature, the unlearning targets were generally confined to the training data, as the model’s memory is learned directly from the training data [cao2015towards, bourtoule2021machine]. We refer to these traditional unlearning tasks as input unlearning, meaning they focus on unlearning the input data used to train the model. A more detailed discussion of different unlearning targets is provided in Section II-A4.
For clarity and simplicity, in input unlearning, we define the unlearning target here as a subset of the training data (i.e., the forget set $\mathcal{D}_{f}\in\mathcal{D}$). The remaining training data is therefore defined as the retain set, where $\mathcal{D}_{f}\cup\mathcal{D}_{r}=\mathcal{D}$. As shown in Figure 2, a straightforward method of unlearning is to retrain the model from scratch using the retain set (i.e., $\theta^{*}=A(\mathcal{D}_{r})$) However, with the increasing size of both data and models in modern applications, retraining from scratch becomes computationally prohibitive and impractical. This challenge has, in turn, driven the development of more efficient methods for machine unlearning. The retrained model $\theta^{*}$ serves as the ground truth for unlearning. Therefore, the goal of unlearning is to design an unlearning algorithm $U(\cdot)$ such that
@@
$\quad\text{s.t.}\quad Cost(U(A,\mathcal{D}_{f},\mathcal{D}_{r},\theta_{0}))<Cost(A(\mathcal{D}_{r})),$
where the input of $U(\cdot)$ may be reduced under different conditions (e.g., when the retain set is not available), $Cost(\cdot)$ denotes the computational overhead, and $\doteq$ denotes the degree of unlearning completeness, i.e., to what extent the target has been completely unlearned. This is also referred to as the effectiveness or efficacy of unlearning in the literature. The discussion of evaluation metrics of unlearning completeness will be provided in Section II-B. For clarity, in this paper, we refer to $\theta$ as the original model, $\theta_{u}=U(A,\mathcal{D}_{f},\mathcal{D}_{r},\theta_{0})$ as the unlearned model, and $\theta^{*}=A(\mathcal{D}_{r})$ as the retrained model.
As illustrated in Figure 2, recommendation models may inadvertently encode sensitive attribute information, constituting a distinct category of data requiring protection beyond training datasets. Crucially, this attribute information is not directly used during model training, meaning that i) it cannot be inferred solely from training data leakage, and ii) it remains unremovable through retraining from scratch (i.e., all input unlearning methods). Consequently, attribute unlearning is fundamentally defined as an attack-based paradigm. A detailed introduction of attribute unlearning and the attack-based evaluation metric are provided in Section III.B and Section IV.B, respectively.
@@ II-A2 Recommender systems
Recommender systems leverage historical interactions to predict user preferences. A fundamental and classical research focus is the rating prediction task, where users rate various items [steck2013evaluation]. Recommendation models leverage observed (i.e., historical) ratings to predict unobserved ones. Collaborative filtering is the foundational approach in modern recommender systems, based on the assumption that users tend to favor similar items, and that an item is likely to be favored by similar users, creating a collaborative effect [koren2021advances]. Under this approach, matrix factorization-based models have become widely used in recommender systems, both in academia and industry [ma2008sorec, hu2008collaborative]. Existing research on recommendation unlearning also primarily focuses on matrix factorization-based models. Generally, the core idea behind matrix factorization-based models is to learn a user embedding matrix and an item embedding matrix. The predicted ratings are then obtained through the dot product (or other combination methods) of these two matrices. Subsequent research has leveraged deep learning techniques [xue2017deep, he2017neural], graph learning approaches [he2020lightgcn, peng2022svd, zheng2021dgcn], and Large Language Models (LLMs) [acharya2023llm, lin2024data] to further enhance recommendation performance.
@@ II-A3 Recommendation Unlearning
  • Last but not least, due to data sparsity and contextual dependencies among users and items, recommendation models are highly sensitive to the quality of the training data [schafer2007collaborative]. This creates a need for systems to unlearn problematic data proactively.
@@
However, traditional machine unlearning methods cannot be indiscriminately applied to recommender systems, due to the unique challenges posed by collaborative interactions and parameters of recommendation models. The collaborative nature between user-item interactions introduces contextual dependencies, meaning a change in one data point can propagate through the system, affecting recommendations for multiple users or items. Firstly, traditional exact unlearning methods often require dividing the dataset. As shown in Figure 3, this additional breakdown of user-item interactions can lead to a decline in recommendation performance. Conversely, the retain set can leverage the collaborative effects to re-remember the forget set, thereby causing non-compliance with the unlearning request. Secondly, approximate unlearning methods typically rely on gradient updates, which do not consider the gradients of associated data samples. As shown in Figure 3, the collaborative effect may cause the model to re-memorize unlearned data through these associated samples. Thirdly, the user and item embeddings, which are crucial parameters in recommendation models, capture the features of users and items. As a result, existing recommendation unlearning methods typically focus on manipulating these embeddings. However, because the user and item embeddings form high-dimensional and dense matrices, and the number of users and items is massive, applying traditional unlearning methods to these embeddings introduces significant computational overhead, which makes recommendation unlearning even more challenging.
II-A4 Unlearning Targets
@@
In addition to the training data, data that does not participate in training may also need to be unlearned in recommender systems. This is because adversaries can potentially infer private information from a trained model, even if that information was never explicitly included in the training data. This type of information, referred to as attributes, is implicitly learned by the model during training. Such attacks are known as attribute inference attacks [gong2018attribute, zhao2021feasibility, jayaraman2022attribute]. The task of unlearning these attributes is called attribute unlearning, which serves as a defense mechanism against attribute inference attacks [ganhor2022unlearning, li2023making, chen2024post]. As shown in Figure 2, the unlearning target of attribute unlearning is the latent user attributes that are not part of the training data, e.g., gender, age, and race.
@@ II-A5 Unlearning Workflow
II-B Design Principles
The goal of unlearning is not merely to eliminate the memory of the target being unlearned. It encompasses broader goals. Generally, there are three key design principles for unlearning methods, which are also applicable to recommendation scenarios [chen2022recommendation, li2024making, li2023ultrare].
@@ II-B Design Principles
  • Algorithmic perspective: Only retraining from scratch (i.e., unlearning from the algorithmic level) satisfies the definition of complete unlearning in this perspective. Thus, this definition of completeness can only be self-evaluated algorithmically or verified by providing training checkpoints [jia2021proof, thudi2022necessity, weng2024proof]. Exact unlearning methods adhere strictly to this definition, designing efficient strategies to achieve retraining. All other unlearning methods are classified as approximate unlearning [nguyen2022survey].
@@
The specific examples of evaluation metrics are introduced in Section IV-B. From the point of unlearning techniques, these four perspectives form a continuum from strict to relaxed definition: beginning with the algorithmic requirement, progressing to parametric similarity, then to functional equivalence, and finally to resilience against adversarial attacks. However, in practical applications, the attack and functional perspectives are particularly significant, as they directly reflect model behavior and security in real-world scenarios. Recent study has demonstrated that satisfying the algorithmic perspective does not necessarily guarantee completeness under the functional or attack perspectives [CURE4Rec2024].
@@ III-A Input Unlearning
In traditional recommendation scenarios, based on the types of models that can be applied, existing input unlearning methods can be categorized into two main approaches: model-agnostic and model-specific. The term model-agnostic refers to methods that can be applied regardless of the model structure, while model-specific refers to methods designed for a particular type of model. In addition to the traditional scenario, we also review the unlearning methods for recommendation tasks in specific scenarios, such as federated learning and LLMs, where the recommendation model differs from traditional recommender systems.
@@ III-A1 Model-agnostic Methods
Model-agnostic unlearning methods for recommendation tasks are largely inspired by traditional machine unlearning techniques used in classification tasks. Based on the definition of unlearning completeness, we further classify the model-agnostic approach into exact unlearning and approximate unlearning [nguyen2022survey].
Figure 7: Different designs for exact unlearning in recommender systems, where $\theta_{0}$, $\theta^{*}$, and $\theta_{u}$ denote the original model, the model retrained from scratch, and the unlearned model respectively, $\mathcal{D}_{i}$ represents the subset of original dataset $\mathcal{D}$, and $\theta_{i}$ represents the corresponding model trained on $i$-th subset. The black and blue arrows depict the workflows of learning and unlearning respectively.
@@
As mentioned in Section II-B, exact unlearning follows a strict definition of unlearning completeness, achieving it at the algorithmic level. Inspired by SISA [bourtoule2021machine], exact recommendation unlearning methods predominantly adopt the ensemble retraining framework. As shown in Figure 7, this framework divides the original dataset into multiple subsets, trains a sub-model on each subset, and aggregates all sub-models into the final model, similar to an ensemble learning system. Note that, to ensure algorithmic unlearning completeness, the design of sub-models is typically identical to that of the original model, including the model structure, hyperparameter settings, and other factors. This framework allows for more efficient retraining during unlearning. When an unlearning request is submitted, the ensemble retraining framework only needs to retrain the subset containing the unlearning targets, avoiding the need to retrain the entire dataset from scratch, thereby enhancing unlearning efficiency.
@@
Li et al. [li2024enhancing] directly apply SISA to recommendation models in intelligence education. Their approach enhances the personalization and accuracy of educational recommendations by selectively forgetting the data inputs for each user.
@@
Building on the design of SISA, Chen et al. propose RecEraser, which introduces two key modifications tailored for recommendation tasks [chen2022recommendation]. First, RecEraser incorporates a balanced clustering module for dataset division, grouping similar users or items into the same subset to preserve the collaborative effects within the data, in contrast to the random division used in SISA. Second, RecEraser adds an attention network to learn the weights for the weighted aggregation of sub-models. This adaptive weighted aggregation, compared to the average weighting or majority voting in SISA, further enhances recommendation performance for the ensemble retraining framework.
@@
Due to the collaborative effect of recommendation data, there is a significant trade-off between unlearning efficiency and model utility. Specifically, increasing the number of dataset divisions can enhance unlearning efficiency, but this also disrupts the collaboration among data, which in turn reduces model utility. To address this issue, Li et al. propose UltraRE, a lightweight modification of RecEraser [li2023ultrare]. UltraRE introduces a new balanced clustering algorithm based on optimal transport theory, which improves both efficiency and clustering performance simultaneously. Additionally, UltraRE simplifies the attention network used during aggregation, replacing it with a logistic regression model to further enhance efficiency.
@@
To further enhance model utility, LASER adopts sequential training during aggregation, rather than parallel training [li2024making]. As shown in Figure 7, sequential training involves training one model on a data subset sequentially. This approach helps mitigate the negative impact of dataset division on collaboration. LASER introduces the concept of curriculum learning to optimize the training sequence of data subsets, thereby improving model utility. However, sequential training also reduces unlearning efficiency. To address this issue, LASER introduces early stopping and parameter manipulation.
However, there are several drawbacks of model agnostic exact unlearning. Note that not all exact unlearning methods rely on ensemble retraining frameworks, nor do they all suffer from the following drawbacks. Some model-specific approaches can achieve high efficiency, e.g., KNN unlearning [schelter2023forget].
@@
  • When samples from the forget set are distributed across multiple subsets, all affected subsets require updates. This creates substantial challenges for subset division design.
  • The efficiency gains from dataset division are limited and incremental, whereas approximate unlearning methods can often provide efficiency improvements that are orders of magnitude better. Additionally, as noted by [li2023ultrare], exact unlearning introduces a trade-off between efficiency and utility, which is particularly significant in recommendation tasks.
  • The core idea of the ensemble retraining framework is trading space for time. Although this framework enhances efficiency, it requires maintaining multiple copies of model parameters, which introduces significant storage overhead. For large-scale recommendation models in practical applications, this creates both storage capacity challenges and potential security concerns.
  • The performance of exact unlearning is suboptimal in practice [CURE4Rec2024]. Although exact unlearning achieves perfect completeness in theory, its empirical performance is unsatisfactory, often worse than approximate unlearning. In real-world scenarios, users prioritize good performance over theoretical guarantees.
Figure 8: Different designs for approximate unlearning in recommender systems, i.e., reverse unlearning and active unlearning. $\theta_{0}$ and $\theta_{u}$ denote the original model and the unlearned model respectively. $\mathcal{D}$ and $\mathcal{D}_{f}$ represents the original datasets and the forget set respectively.
@@
Reverse Unlearning. Deep learning models are typically trained using gradient descent-based optimization. An intuitive approach for unlearning is to add back the gradient of the target data that was previously subtracted, thereby mitigating its influence on the model and achieving the goal of unlearning from a parametric perspective. Formally, let us assume that $z_{i}\in\mathcal{D}$ is a data record from the training set, and the recommendation loss function is $\ell(z_{i},\theta)$. We further assume that the loss function is twice-differentiable and strictly convex, which holds true for the majority of recommendation loss functions. Note that these assumptions serve only as foundational simplifications for reduction-based analyses and do not imply convexity in deep learning-based recommendation models. Following the empirical risk minimization framework, the optimal model parameter is given by
$\theta_{0}=\underset{\theta}{\operatorname{\arg\min}}\hskip 2.84526pt\sum_{i=1}^{n}\ell(z_{i},\theta),$ (2)
@@
$\theta^{*}=\underset{\theta}{\operatorname{\arg\min}}\hskip 2.84526pt\sum_{i=1}^{n}\ell(z_{i},\theta)-\ell(z,\theta).$ (3)
The early study in traditional machine unlearning explores directly adding back the first-order gradient [graves2021amnesiac]. However, due to the stochastic nature and randomness inherent in deep learning, the first-order gradient is not always accurate for estimating the influence of the target data. Consequently, subsequent research has focused on more precisely estimating the influence and directly subtracting this influence from the original model [sekhari2021remember]. This process is simply formulated as: